OpenStudy (anonymous):

can somebody plz help me understand this?? log1/5 625

OpenStudy (saifoo.khan):

simply use calc.

OpenStudy (anonymous):

idk how 2 plug it in

OpenStudy (anonymous):

i wanna learn how to do it by my self plz help

OpenStudy (saifoo.khan):

\[\Huge \log \frac{1}{5}625 \]

OpenStudy (anonymous):

Did youmean? $$ log_{1/5} 625$$

OpenStudy (saifoo.khan):

is it this?

OpenStudy (anonymous):

then it is -4

OpenStudy (anonymous):

yes, how did you do it /

OpenStudy (saifoo.khan):

who you?

OpenStudy (anonymous):

By the definition of logarithm $$log_ab=c <=> a^c=b $$ In your case a=1/5 b=625 $$(1/5)^{-4}=5^4=625$$ So -4 is the right answer. There is no way to calculate logarithms exept guessing or using calculators(or Taylor series, but then the value will not be exact)